OpenStudy (anonymous):

Medal Chance!! Can anyone convert this answer into a fraction? I don't have the proper calculator to do this. I have the answer...just need to convert it.

OpenStudy (anonymous):

3.2_{E} -4

OpenStudy (mathmate):

E-4 means \(10^{-4}\) For example, 5.4E-6=\(5.4\times 10^{-6}=\dfrac{5.4}{1000000}=\dfrac{54}{10000000}=\dfrac{27}{5000000}\)

OpenStudy (anonymous):

Well...then...Is one of those supposed to be my answer?? I'm taking Pre-Cal right now...

OpenStudy (anonymous):

Would you mind helping me with the question??

OpenStudy (anonymous):

What is the 10th term of the sequence 625, 125, 25,...? * 1 / 3,125 * 1 / 625 * 1 / 1,250 * 1 / 15,625

OpenStudy (mathmate):

Please, do not expect a direct answer from the responses. This will deprive you from learning what is needed for your studies (and exams). Read and understand the help given to you, and apply to your problem. Give the problem a try, and if you still have difficulties, show what you did and where you got stuck. :)

OpenStudy (anonymous):

I'm not asking for the answer. I'm asking for help on how to solve it. I'm really stuck, and the answer lies between these 4. Please help.

OpenStudy (mathmate):

If you follow the example using the numbers from the question and you should not be far from the answer.